OpenStudy (anonymous):

solve x^3 - 3x + 5 using the definition of the derivative

OpenStudy (thomas5267):

Differentiate or solving this equation?

OpenStudy (anonymous):

differentiate

OpenStudy (thomas5267):

Definition of derivatives: \[ f'(x)=\lim_{h\to0}\frac{f(x+h)-f(x)}{h} \]

OpenStudy (anonymous):

yes but when i plug it in then the answer becomes complicated i get to

OpenStudy (anonymous):

\[[(x+h)^3 - 3(x+h) + 5 * x^3 -3x +5]/h \]

OpenStudy (thomas5267):

\[ \lim_{h\to 0}\frac{f(x+h)-f(x)}{h},\text{ not}\lim_{h\to 0}\frac{f(x+h)\times f(x)}{h} \]

OpenStudy (anonymous):

oh! let me retry it!

OpenStudy (anonymous):

i get to the point where i do (x+h)^3 and i dont really know what it comes out to i end up with x^3 + 2x^2h + h^2x + hx^2 + 2xh^2 + h^3 and i dont know how to simplify this

OpenStudy (thomas5267):

You can directly use binomial theorem to expand \((x+h)^3\). Binomial theorem: \[ (x+y)^n=\sum_{k=0}^n\binom{n}{k}x^ky^{n-k} \] I don't know how to explain the theorem though.

OpenStudy (anonymous):

i dont understand that theorem either

OpenStudy (luigi0210):

There is another way, since it's a small power you could use pascal's triangle.

OpenStudy (luigi0210):

Just to get the constants that is.

OpenStudy (larseighner):

With lower powers, multiplying longhand is often the fastest, simplest way.

OpenStudy (larseighner):

This what longhand multiplication looks like \(\large \displaystyle \begin{align} x &+ h \cr x &+ h \cr \hline hx &+ h^2 \cr x^2 + hx& \cr\hline x^2 + 2h&x + h^2 \cr &x + h \cr \hline hx^2 +2h^2&x + h^3 \cr x^3 +2hx^2 +h^2x& \cr \hline x^3 + 3hx^2 +3h^2&x + h^3 \end{align}\)

OpenStudy (anonymous):

thank you for explaining it clearly. i understand how to do it now thanks everyone